# SEMPO - Retrieving poles, residues and zeros in the complex frequency plane from an arbitrary spectral response [^1] This work was funded by the French National Research Agency ANR Project DILEMMA (ANR-20-CE09-0027)

### v1

@@ Abstract
The Singularity Expansion Method Parameter Optimizer - SEMPO - is a toolbox to extract the complex poles, zeros and residues of an arbitrary response function acquired along the real frequency axis. SEMPO allows to determine this full set of complex parameters of linear physical systems from their spectral responses only, without prior information about the system. The method leverages on the Singularity Expansion Method of the physical signal. This analytical expansion of the meromorphic function in the complex frequency plane motivates the use of the Cauchy method and auto-differentiation-based optimization approach to retrieve the complex poles, zeros and residues from the knowledge of the spectrum over a finite and real spectral range. Both approaches can be sequentially associated to provide highly accurate reconstructions of physical signals in large spectral windows. The performances of SEMPO are assessed and analysed in several configurations that include the dielectric permittivity of materials and the optical response spectra of various optical metasurfaces.
@@ 1 Introduction
The spectral response of a linear physical system carries rich information on the system and on its interaction with excitation fields. The complex frequency space is well suited to perform this analysis, and the poles of the different transfer operators or scattering channels provide much of this information on the system. Different methods have been developed in order to model the response of the system to a given excitation with parameters determined in the complex frequency space including the Singularity Expansion Method (SEM), the Resonant State Expansion, or the quasi-normal modes theory [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
@@
The SEM provides an expansion of any physical signal, such as the elements of a scattering operator, in terms of its poles and residues or alternatively in terms of its poles and zeros determined in the complex frequency plane. The expressions resulting from the distributions of log-singularities (poles and zeros) and residues can be used to approximate functions at any frequency over large spectral windows, with small and discrete sets of parameters [11, 12, 13, 14, 15]. They provide simple models to describe the behaviour of scatterers and cavities [16, 4, 17, 18]. The behaviour of physical systems and their interaction with excitation signals can be explained in terms of poles and zeros [19, 20, 21, 22, 23, 24].
@@
Information about the distribution in the complex frequency plane being encoded in the spectral response of the physical system, the challenge is therefore to decode this information. Different methods have been proposed to retrieve the log-singularities and residues from a spectral response either obtained in the complex $\omega$ plane or along the real frequency axis. When the available data are acquired or generated in the complex frequency plane, the calculation of the poles and residues (or equivalently the poles and zeros) is quite straightforward through contour integrals [25, 26, 22, 27]. However, in practice, many spectra are acquired over a real spectral range $[\omega_{A},\omega_{B}]\subset\mathbb{R}$. In the latter case, much fewer methods exist but this longstanding problem has been attracting more and more attention [28, 29, 30, 31, 32]. Existing methods mainly approximate the response functions as meromorphic functions, the parameters of which are either the poles and zeros (Cauchy method [31] and AAA algorithm [32] indirectly), or the poles and residues (matrix pencil method [28], harmonic inversion [29], vector fitting [30]). They rely on matrix algebra, but can be quite sensitive to an initial guess of some parameters, as in the case of the vector fitting method.
@@
Here, we describe two approaches to identify, in the complex frequency plane, the log-singularities (poles and zeros) and residues of a physical system from spectra acquired over a finite range of real frequencies. The first approach is based on the Cauchy method [14, 31, 33] and works by expressing the function through a rational approximation. In this work, we describe the working principle of the original method and explain how it can be improved and optimized using the properties of physical systems. The second method relies on auto-differentiation [34, 35, 36, 37, 38, 24] to solve optimization problems through gradient-descent. The formulation of these optimization problems is detailed, as well as their application to retrieve the SEM.
@@
The physical signal that we consider is a transfer function such as an element of the scattering $S$-matrix that is represented by a meromorphic function $h(\omega)$ which can be expanded using the SEM [1, 39, 4, 40, 5], i.e.
@@
with $z^{(\ell)}$ the zeros of $h(\omega)$, and $\eta_{0}$ a known constant which depends on $z^{(\ell)}$, $p^{(\ell)}$, and the response $h(a)$ to an arbitrary frequency $a$.
The distribution of poles $p^{(\ell)}$, zeros $z^{(\ell)}$, and residues $r^{(\ell)}$ fully characterize the response function $h(\omega)$. Therefore, the log-singularities are intrinsic properties of physical systems, and their retrieval is relevant. However, there is a potentially infinite number of parameters, which implies that they are never all taken into account since the SZF and SEM must be numerically truncated. In the following we describe how to retrieve the poles, residues and zeros once we perform this truncature. Furthermore, we show how to combine both methods to improve the overall performance.
@@ 2.1 Overview of the original method
Let us first present an overview of the Cauchy method to retrieve the poles and zeros of a function from a set of measurements acquired along a real finite spectral range.
@@
$[\mathbf{A},-\mathbf{B}]\begin{bmatrix}\mathbf{a}\\ \mathbf{b}\\ \end{bmatrix}=\mathbf{0}.$ (8)
@@
$\mathbf{A}=\begin{bmatrix}1&\omega_{1}&...&\omega_{1}^{M_{z}}\\ 1&\omega_{2}&...&\omega_{2}^{M_{z}}\\ &...&\\ 1&\omega_{N}&...&\omega_{N}^{M_{z}}\\ \end{bmatrix}\$ (9)
@@
The Cauchy method constrains the kernel of the matrix $\mathbf{C}=[\mathbf{A},~{}-\mathbf{B}]$ in order to obtain a unique algebraic solution up to a scaling factor. The number of unique frequencies $N$ is usually greater than the number of columns $2+M_{p}+M_{z}$ of $\mathbf{C}$. It follows that the rank $r$ of $\mathbf{C}$ is given by
@@
The vector $\mathbf{b}$ is thus set to the eigenvector associated with the 00 eigenvalue of $\mathbf{R}_{22}$, which is uniquely defined up to a multiplicative constant since $K=1$. We can then obtain $\mathbf{a}$ from Equation (13):
@@
$e^{(2)}_{r,M_{p},M_{z}}\left[\hat{\mathbf{h}},\mathbf{h}\right]=\frac{{||\hat{\mathbf{h}}-\mathbf{h}||}_{2}}{{||\mathbf{h}||_{2}}}$ (16)
@@ 2.2 Accuracy-driven Cauchy method
where $\mathbf{U}$ and $\mathbf{V}$ are two unitary matrices, $\mathbf{V}^{H}$ is the conjugate transpose of $\mathbf{V}$, $r=M_{z}+M_{p}+2$, and $\sigma_{1},\sigma_{2},...,\sigma_{r_{+}1}$ are the singular values of $\mathbf{C}_{\lambda}$. By setting $\sigma_{r+1}$ to 00, we enforce $r$ as the rank of the resulting matrix $\hat{\mathbf{C}}_{\lambda}$, ensuring a dimension of the kernel $K=1$. We then perform the classical Cauchy method on all matrices $\hat{\mathbf{C}}_{\lambda}$, resulting in $\Lambda$ solutions $[\mathbf{a}_{\lambda},\mathbf{b}_{\lambda}]$, each associated with a reconstruction $\hat{\mathbf{h}}_{\lambda}$ using Equations (3) to  (5).
@@ 2.3 Physics-informed corrections
If $h(\omega)$ is the response or the transfer function of a physical system, it is Hermitian-symmetric, i.e. its poles and zeros come in pairs $(p^{(\ell)},-\overline{p^{(\ell)}})$ and $(z^{(\ell)},-\overline{z^{(\ell)}})$, and the constant $\eta_{0}$ is real-valued. We can rewrite the truncated SZF [5] as
@@
where $\mathrm{i}z_{\mathrm{i}\mathbb{R}}$ are purely imaginary zeros, $-\mathrm{i}p_{\mathrm{i}\mathbb{R}}$ are purely imaginary poles, $z_{\mathbb{C}}$ are non-imaginary zeros, and $p_{\mathbb{C}}$ are non-imaginary poles. The number of poles and zeros read as $M_{z}=M_{z,\mathrm{i}\mathbb{R}}+M_{z,\mathbb{C}}$ and $M_{p}=M_{p,\mathrm{i}\mathbb{R}}+M_{p,\mathbb{C}}$.
In order to take into account this Hermitian symmetry, we add negative frequencies, effectively multiplying the number of frequencies by 2. For every frequency $\omega_{i}$ associated with the value $h_{i}$, the opposite frequency $-\omega_{i}$ is added, along with $\overline{h_{i}}$. This operation is performed at the beginning of the Cauchy method, right before the step corresponding to Equation (6).
@@
Finally, we propose to further modify the distribution of poles to account for the usually assumed stability of physical systems or signals. Stable poles possess a negative imaginary part. Therefore, we penalize the solutions associated with unstable poles in the ADC method by multiplying the errors $e^{(2)}_{r,M_{p},M_{z}}\left[\mathbf{h}_{\lambda},\mathbf{h}\right]$ by $(1+M_{inst,p})$ with $M_{inst,p}$ the number of poles with a strictly positive imaginary part. What is more, we add a negative imaginary part $-\mathrm{i}q_{0}$ to the poles located on the real frequency axis once the optimal solution is obtained. $q_{0}$ is yet another free parameter set to $10^{-5}$ by default. Since the effect of this modification on the zeros is not easily obtained, this last step is performed after converting the SZF into the SEM, i.e. by considering the poles and their associated residues instead of considering the poles and zeros.
@@ 3.1 Formulating the optimization problems
While it is possible to deal with complex values in the optimization process, current auto-differentiation tools have only recently been able to deal with such numbers, and some functions are still not properly defined in $\mathbb{C}$ [42, 43]. We thus express the parameters in $\mathbb{R}^{2}$ instead of $\mathbb{C}$. We split the $M$ poles of Equation (26) into $M_{I}$ purely imaginary poles $\mathrm{i}q^{(\ell)}$ with residues $\mathrm{i}\rho^{(\ell)}$ and $M_{C}$ complex pairs of poles $p^{(\ell)}=p_{R}^{(\ell)}+\mathrm{i}p_{I}^{(\ell)}$ with residues $r^{(\ell)}=a^{(\ell)}+\mathrm{i}b^{(\ell)}$.
We obtain a form of the SEM expression that is compatible with the auto-differentiation tools [24]:
$h(\omega)=h_{NR}+\sum_{\ell=1}^{M_{I}}\frac{\rho^{(\ell)}}{q^{(\ell)}}+\sum_{\ell=1}^{M_{C}}\left[\frac{a^{(\ell)}+\mathrm{i}b^{(\ell)}}{\omega-p_{R}^{(\ell)}-\mathrm{i}p_{I}^{(\ell)}}-\frac{a^{(\ell)}-\mathrm{i}b^{(\ell)}}{\omega+p_{R}^{(\ell)}-\mathrm{i}p_{I}^{(\ell)}}\right].$ (27)
The non-resonant term is also a real-valued parameter due to the Hermitian symmetry. The explicit discrimination between the $M_{I}$ imaginary poles and the $M_{C}$ pairs of poles is not necessary, but it can help in improving the results if imaginary poles are expected. In addition, it reduces the number of parameters from $1+4M$ to $1+2M_{I}+4M_{C}$ which can slightly improve the speed of the optimization process.
We wish to find a set of parameters $\mathcal{P}$ defined as
$\mathcal{P}=\left\{h_{NR},~{}\left(\rho^{(\ell)}\right)_{\ell},\left(q^{(\ell)}\right)_{\ell}\left(a^{(\ell)}\right)_{\ell},\left(b^{(\ell)}\right)_{\ell},\left(p_{R}^{(\ell)}\right)_{\ell},\left(p_{I}^{(\ell)}\right)_{\ell}\right\}$ (28)
for which a small value of the arbitrary loss function $\mathcal{L}$ is reached. Ideally, we want to retrieve the optimal set of parameters $\mathcal{P}^{*}$ defined as
$\mathcal{P}^{*}=\operatorname*{arg\,min}_{\mathcal{P}}\mathcal{L}[\mathbf{h},\hat{\mathbf{h}},\mathcal{P}]$ (29)
where $\mathcal{L}[h_{i},\hat{h}_{i},\mathcal{P}]$ evaluates the distance of $\hat{h}_{i}$ to $h_{i}$ using the parameter set $\mathcal{P}$ in the truncated SEM from Equation (26) to reconstruct the function. The minimization of the loss function is performed via a gradient-descent(-like) method. Classically, the loss function is limited to the set of functions for which this gradient can be computed.
Nevertheless, the recent development of auto-differentiation (for machine learning in particular), has provided means to explore more exotic loss functions. The tools provided in the PyTorch library allow to freely design the loss function while the gradient is automatically calculated by the library through the chain rule.
The loss function that we use requires the definition of the relative $L_{2}$ and $L_{\infty}$ errors:
$e_{q}\left[\mathbf{h},\hat{\mathbf{h}}\right]=\frac{{||\mathbf{h}-\hat{\mathbf{h}}||}_{q}}{{||\mathbf{h}||_{q}}}$ (30)
@@
$\begin{gathered}{||\mathbf{v}||}_{2}=\sqrt{\sum_{i=1}^{N}|v_{i}|^{2}}\\ {||\mathbf{v}||}_{\infty}=\max_{i}|v_{i}|.\end{gathered}$ (31)
@@
$\mathcal{L}[\mathbf{h},\hat{\mathbf{h}},\mathcal{P}]=\alpha_{1}e^{(2)}\left[\mathbf{h},\hat{\mathbf{h}}\right]+\alpha_{2}\left\lVert\frac{\mathbf{h}-\hat{\mathbf{h}}}{\mathbf{h}}\right\rVert_{\infty}+\alpha_{3}\left<\frac{\delta\mathbf{h}_{R}}{\mathbf{h}_{R}}\right>+\alpha_{4}\left<\frac{\delta\mathbf{h}_{I}}{\mathbf{h}_{I}}\right>$ (32)
where the notation <.><.>< . > corresponds to the averaging operation over all the sampled frequencies. $\delta\mathbf{h}_{R}$ is the vector of errors between the real part of the model $\hat{\mathbf{h}}$ and the experimental data $\mathbf{h}$, and $\mathbf{h}_{R}$ is the vector of the absolute values of the real part $\mathbf{h}$. It is used to change the weight given to the frequency $\omega_{i}$ in $\mathbf{\delta h_{R}}$. $\mathbf{\delta h_{I}}$ and $\mathbf{h}_{I}$ do the same for the imaginary part:
[…]
@@
In an accuracy-driven optimization, these parameters, along with the numbers of poles $M_{I}$ and $M_{C}$, can be optimized by grid-search as to validate a criterion such as the minimization of the relative $L_{2}$ error between $h_{i}$ and $\hat{h}_{i}$.
However, this could lead to a fitting of the noise rather than fitting the signal brought about by the data. We have found that setting $\alpha_{1}=1$, and keeping the other weights small often provides good results. In the following examples, the values of $\mathbf{\alpha}$ are therefore intuited based on the expected response curve.
Figure 4: (a) 2D array of square Ag pillars of width $w=85$ nm and height $h=35$ nm, and period $p=130$ nm over a substrate of Ag. (b) The signal is the 00-th order reflection coefficient $h(\omega)=r_{00}(\omega)$ of the 2D array depicted in (a) illuminated in normal incidence reconstructed with the SEM in Equation (27) (red curve), and compared to simulated data (blue markers). The absolute error $|h_{i}-\hat{h}_{i}|$ (black curve) is less than $10^{-3}$ at all frequencies, and a relative $L_{2}$ error of $4.32\times 10^{-3}$ is obtained. (c) Distribution of poles in the complex $\omega$ plane.
Using the example of the 00-th order reflection coefficient of a 2D grating of silver described in Figure 4 (a), we show the result of the retrieval of the singularity expansion of Equation (27) in Figure 4 (b,c), with $M_{I}=0$ imaginary poles, $M_{C}=7$ complex pairs of poles, and $\mathbf{\alpha}=(1,0,0.2,0.2)$.
These weight coefficients are such that the main focus is put on the relative $L_{2}$ error, with a smaller contribution of the errors on the real and imaginary parts, individually. We chose them after observing variations of equal amplitude of the real and imaginary parts in the target function. The non-imaginary poles were initialized with their real parts evenly distributed from $2.5$ to $4.9\times 10^{15}$ rad/s, and their imaginary part set to $-0.05$ times their real parts. We obtain a relative $L_{2}$ error of $4.32\times 10^{-3}$%, with a Hermitian-symmetric solution.

### v2

# SEMPO - Retrieving complex poles, residues and zeros from arbitrary real spectral responses [^1] This work was funded by the French National Research Agency ANR Project DILEMMA (ANR-20-CE09-0027)

@@ Abstract
The Singularity Expansion Method Parameter Optimizer - SEMPO - is a toolbox to extract the complex poles, zeros and residues of an arbitrary response function acquired along the real frequency axis. SEMPO allows to determine this full set of complex parameters of linear physical systems from their spectral responses only, without prior information about the system. The method leverages on the Singularity Expansion Method of the physical signal. This analytical expansion of the meromorphic function in the complex frequency plane motivates the use of an accuracy-driven improved version of the Cauchy method constrained by properties of physical systems, as well as an auto-differentiation-based optimization approach. Both approaches can be sequentially associated to provide highly accurate reconstructions of physical signals in large spectral windows. The performances of SEMPO are assessed and analysed in several configurations that include the dielectric permittivity of materials and the optical response spectra of various optical metasurfaces. SEMPO’s performances are thoroughly analyzed and benchmarked with other state-of-the-art methods to highlight its capability to retrieve the natural poles of a physical system.
Program summary
  • Program title: Singularity Expansion Method Parameter Optimizer (SEMPO)
  • CPC Library link to program files:
  • Developer’s repository link: https://doi.org/10.5281/zenodo.15210008
  • Licensing provisions: Creative Commons Attribution 4.0 International
  • Programming language: Python
  • Nature of problem: Spectral functions of scattering coefficients of physical systems can be rigorously expanded through the Singularity Expansion Method. This expansion method permits to cast the spectral function through two equivalent forms: a sum of complex rational functions depending on complex poles and residues, and a factorized expression that depends on complex poles and zeros. Spectral functions are usually acquired along the real frequency axis, the challenge is thus to identify the set of complex poles, zeros and residues from these acquisitions at real frequencies.
  • Solution method: SEMPO is a numerical toolbox that aims at extracting the complex poles, zeros and residues from arbitrary spectral functions. SEMPO relies on both an improved algebraic Cauchy method associated with the factorized expression, and an open-source auto-differentiation method associated with the expansion in terms of poles and residues. Both methods are based on the Singularity expansion method and leverage the properties of physical systems and signals. The accuracy of SEMPO is thoroughly assessed and analysed in different configurations including the spectral responses of optical metasurfaces and the dielectric permittivity of optical materials.
@@ 1 Introduction
The spectral response of a linear physical system carries rich information on the system and on its interaction with excitation fields. The complex frequency space is well suited to perform this analysis, and the poles and zeros of the different transfer operators or scattering channels provide much of this information on the system. Different methods have been developed in order to model the response of the system to a given excitation with parameters determined in the complex frequency space including the Singularity Expansion Method (SEM), the Resonant State Expansion, or the quasi-normal modes theory [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The development of these methods has been motivated by the need to tailor and optimize resonances that are cumbersome in wave physics; Resonances are at the core of cavity-driven systems in water waves [12, 13, 14], acoustics [15], or cavity opto-mechanics [16, 17]. They play a key role in optics and photonics in enhancing light-matter interactions [18, 19, 20, 21, 22, 23], cavity electrodynamics [24], or designing highly efficient laser cavities [25]. The strength of resonances is characterized by the quality factor that can be calculated through several methods. It can be first estimated under the assumption of well-defined Lorentzian spectral responses by taking the ratio between the central resonant frequency and the full width at half maximum of the resonant response [26]. It can also be calculated with the temporal coupled mode theory through the calculation of the radiative and internal loss times [27, 28] or in the harmonic regime by calculating the ratio between the real and the imaginary part of a complex singularity [29, 30]. Recently, it was showed that the concept of quality factor can be generalized by a meromorphic quality function proportional to the Wigner-Smith time delay [31]. An SEM expansion of this function shows that both poles and zeros must be taken into account to characterize resonances. This function identifies with the quality factor at frequencies equal to the real frequency of the singularity. These results illustrate the importance of retrieving poles, zeros and residues in the complex frequency plane.
@@
The SEM provides an expansion of any physical signal, such as the elements of a scattering operator, in terms of its poles and residues or alternatively in terms of its poles and zeros determined in the complex frequency plane. The expressions resulting from the distributions of poles, zeros and residues can be used to approximate functions at any frequency over large spectral windows, with small and discrete sets of parameters [32, 33, 34, 35, 36, 37, 38]. They provide simple models to describe the behaviour of scatterers and cavities [39, 4, 40, 41]. The behaviour of physical systems and their interaction with excitation signals can be explained in terms of poles and zeros [42, 43, 44, 45, 46, 47].
@@
If the complex frequency analysis turns out to be highly relevant for analyzing physical spectra and for characterizing interactions of waves with physical systems, data are in most cases experimentally acquired along the real frequency axis. The challenge is therefore to extract complex parameters from signals acquired along real frequencies. The SEM shows that any linear physical spectral response can be expanded onto an infinite set of complex poles, zeros and residues. It turns out that the distribution of poles, zeros and residues in the complex frequency plane determines the spectral shape of the response function over the real frequencies. This means that the information about their complex distribution is encoded in the spectral response on the real axis, and the challenge is therefore to extract this information from functions of real frequencies. Different methods have been proposed to retrieve the poles, zeros and residues from a spectral response either obtained in the complex $\omega$ plane or along the real frequency axis. When the available data are acquired or generated in the complex frequency plane, the calculation of the poles and residues (or equivalently the poles and zeros) is quite straightforward through contour integrals [48, 49, 45, 50]. However, in practice, many spectra are acquired over a real spectral range $\omega_{A},\omega_{B}]\subset\mathbb{R}$. In the latter case, much fewer methods exist but this longstanding problem has been attracting more and more attention [51, 52, 53, 54, 12, 55, 37, 38, 56]. Existing methods mainly approximate the response functions as meromorphic functions, the parameters of which are either the poles and zeros (Cauchy method [54] and AAA-Adaptive Antoulas-Anderson method- [55] indirectly), or the poles and residues (matrix pencil method [51], harmonic inversion [52], phase analysis [57], vector fitting [53]). They rely on matrix algebra, but can be quite sensitive to an initial guess of some parameters, as in the case of the vector fitting method.
@@
Here, we describe two approaches to identify, in the complex frequency plane, the poles, zeros and residues of a physical system from spectra acquired over a finite range of real frequencies. The first approach is based on the Cauchy method [35, 54, 58] and works by expressing the function through a rational approximation. In this work, we describe the working principle of the original method and explain how it can be improved and optimized as to first minimize the error with a target signal sample in Section 2.1, and then to leverage the properties of physical systems to yield physics-informed corrections in Section 2.2. The second method relies on auto-differentiation [59, 60, 61, 62, 63, 47] to solve optimization problems through gradient-descent. The formulation of these optimization problems is detailed, as well as their application to retrieve the SEM. We apply and analyze the performances of these methods by studying the specular optical response of metasurfaces consisting of periodic nano-elements composed of dispersive materials. We consider 2D arrays of silicon nanodisks on a glass substrate and silver square nanopillars on a silver substrate, and a 1D gold grating consisting of nanoslits etched on a gold substrate. All photonic nanostructures exhibit rich optical spectra useful to assess the performances of SEMPO. Optical responses of these periodic photonic nanostructures are calculated with the Rigorous Coupled Wave Analysis implemented with the software Planopsim and the simulated data are provided in the repository link. Another category of optical response is considered by studying the dielectric permittivity of gold.
@@
The physical signal that we consider is a transfer function such as an element of the scattering $S$-matrix [64, 65]. Each element of the $S$-matrix is represented by a meromorphic function $h(\omega)$ which can be expanded using the SEM [1, 66, 4, 67, 41, 5], i.e.
@@
with $z^{(\ell)}$ the zeros of $h(\omega)$, and $\eta_{0}$ a known constant which depends on $z^{(\ell)}$, $p^{(\ell)}$, and the response $h(a)$ to an arbitrary frequency $a$. Let us point out that in lossless stable systems, i.e. when the amplitude of $h(\omega)$ is constant over the real frequency axis, as is the case as a first approximation with the reflection coefficient of a perfect mirror for instance, then the poles and zeros are complex conjugates. In this case, the SZF identifies with the Blaschke product [69].
The distribution of poles $p^{(\ell)}$, zeros $z^{(\ell)}$, and residues $r^{(\ell)}$ fully characterize the response function $h(\omega)$. Therefore, the poles and zeros are intrinsic properties of physical systems, and their retrieval is relevant. However, there is a potentially infinite number of parameters, which implies that they are never all taken into account since the SZF and SEM must be numerically truncated. In the following we describe how to retrieve the poles, residues and zeros once we perform this truncature. Furthermore, we show how to combine both methods to improve the overall performance.
@@ 2.1 Overview of the original method
Let us first present an overview of the Cauchy method [35, 54, 58] to retrieve the poles and zeros of a function from a set of measurements acquired along a real finite spectral range.
@@
$\mathbf{A},-\mathbf{B}]\begin{bmatrix}\mathbf{a}\\ \mathbf{b}\\ \end{bmatrix}=\mathbf{0}.$ (8)
@@
$\mathbf{A}=\begin{bmatrix}1&\omega_{1}&...&\omega_{1}^{M_{z}}\\ 1&\omega_{2}&...&\omega_{2}^{M_{z}}\\ &...&\\ 1&\omega_{N}&...&\omega_{N}^{M_{z}}\\ \end{bmatrix}\\$ (9)
@@
While the polynomial coefficients $\mathbf{a}$ and $\mathbf{b}$ could be obtained by solving Equation (8) through the least-square method, the Cauchy method instead constrains the kernel of the matrix $\mathbf{C}=[\mathbf{A},~-\mathbf{B}]$ in order to obtain a unique algebraic solution up to a scaling factor. The number of unique frequencies $N$ is usually greater than the number of columns $2+M_{p}+M_{z}$ of $\mathbf{C}$. It follows that the rank $r$ of $\mathbf{C}$ is given by
@@
The vector $\mathbf{b}$ is thus set to the eigenvector associated with the 0 eigenvalue of $\mathbf{R}_{22}$, which is uniquely defined up to a multiplicative constant since $K=1$. We can then obtain $\mathbf{a}$ from Equation (13):
@@
[…]
@@ 2.2 Accuracy-driven Cauchy method
where $\mathbf{U}$ and $\mathbf{V}$ are two unitary matrices, $\mathbf{V}^{H}$ is the conjugate transpose of $\mathbf{V}$, $r=M_{z}+M_{p}+2$, and $\sigma_{1},\sigma_{2},...,\sigma_{r_{+}1}$ are the singular values of $\mathbf{C}_{\lambda}$. By setting $\sigma_{r+1}$ to 0, we enforce $r$ as the rank of the resulting matrix $\hat{\mathbf{C}}_{\lambda}$, ensuring a dimension of the kernel $K=1$. We then perform the classical Cauchy method on all matrices $\hat{\mathbf{C}}_{\lambda}$, resulting in $\Lambda$ solutions $\mathbf{a}_{\lambda},\mathbf{b}_{\lambda}]$, each associated with a reconstruction $\hat{\mathbf{h}}_{\lambda}$ using Equations (3) to  (5).
@@ 2.3 Physics-informed corrections
If $h(\omega)$ is the response or the transfer function of a physical system, we generally assume that it is the Fourier-transform of a real-valued function in the temporal domain. As a consequence, $h(\omega)$ is Hermitian-symmetric, i.e. $h(\omega)=\overline{h(-\overline{\omega})}$. Its poles and zeros come in pairs $(p^{(\ell)},-\overline{p^{(\ell)}})$ and $(z^{(\ell)},-\overline{z^{(\ell)}})$, and the constant $\eta_{0}$ is real-valued [5]. We can rewrite the truncated SZF as
@@
where $\mathrm{i}z_{\mathrm{i}\mathbb{R}}$ are purely imaginary zeros, $\mathrm{i}p_{\mathrm{i}\mathbb{R}}$ are purely imaginary poles, $z_{\mathbb{C}}$ are non-imaginary zeros, and $p_{\mathbb{C}}$ are non-imaginary poles. The number of poles and zeros read as $M_{z}=M_{z,\mathrm{i}\mathbb{R}}+M_{z,\mathbb{C}}$ and $M_{p}=M_{p,\mathrm{i}\mathbb{R}}+M_{p,\mathbb{C}}$.
In order to take into account this Hermitian symmetry, we add negative frequencies, effectively multiplying the number of frequencies by 2. For every frequency $\omega_{i}$ associated with the value $h_{i}$, the opposite frequency $\omega_{i}$ is added, along with $\overline{h_{i}}$. This operation is performed at the beginning of the Cauchy method, right before the step corresponding to Equation (6).
@@
Finally, we propose to further modify the distribution of poles to account for the usually assumed stability of physical systems or signals. When an inverse Fourier-transform is performed to return to the temporal domain, each pole $p^{(\ell)}$ with residue $r^{(\ell)}$ gives rise to a complex exponential term $\mathrm{i}e^{-\mathrm{i}{\mathrm{Re}}[p^{(\ell)}]t}e^{+{\mathrm{Im}}[p^{(\ell)}]t}$, which grows to $\infty$ in modulus if ${\mathrm{Im}}[p^{(\ell)}]>0$ [4, 5]. A stable or non-diverging system or signal must thus possess poles with a negative imaginary part. By extension, a pole is called stable if its imaginary part is negative. To favour stable solutions, we penalize the solutions associated with unstable poles in the ADC method by multiplying the errors $e^{(2)}_{r,M_{p},M_{z}}\left[\mathbf{h}_{\lambda},\mathbf{h}\right]$ by $(1+M_{inst,p})$ with $M_{inst,p}$ the number of poles with a strictly positive imaginary part. What is more, we add a negative imaginary part $\mathrm{i}q_{0}$ to the poles located on the real frequency axis once the optimal solution is obtained. $q_{0}$ is yet another free parameter set to $10^{-5}$ by default. Since the effect of this modification on the zeros is not easily obtained, this last step is performed after converting the SZF into the SEM, i.e. by considering the poles and their associated residues instead of considering the poles and zeros.
@@ 3.1 Formulating the optimization problems
While it is possible to deal with complex values in the optimization process, current auto-differentiation tools have only recently been able to deal with such numbers, and some functions are still not properly defined in $\mathbb{C}$ [70, 71]. We thus express the parameters in $\mathbb{R}^{2}$ instead of $\mathbb{C}$.
The parameters, i.e. the real and imaginary parts of the poles $p^{(\ell)}$ and residues $r^{(\ell)}$, and the real-valued non-resonant term $h_{NR}$, are obtained by minimizing a loss function via a gradient-descent(-like) method. The loss function that we use requires the definition of the relative $L_{2}$ and $L_{\infty}$ errors:
[…]
@@
[…]
@@
$\mathcal{L}[\mathbf{h},\hat{\mathbf{h}}]=\alpha_{1}e^{(2)}\left[\mathbf{h},\hat{\mathbf{h}}\right]+\alpha_{2}\left\lVert\frac{\mathbf{h}-\hat{\mathbf{h}}}{\mathbf{h}}\right\rVert_{\infty}+\alpha_{3}\left<\frac{\delta\mathbf{h}_{R}}{\mathbf{h}_{R}}\right>+\alpha_{4}\left<\frac{\delta\mathbf{h}_{I}}{\mathbf{h}_{I}}\right$> (29)
where $\mathbf{h}$ is the target, $\hat{\mathbf{h}}$ is the prediction, and the notation <.><.> corresponds to the averaging operation over all the sampled frequencies. $\delta\mathbf{h}_{R}$ is the vector of errors between the real part of the model $\hat{\mathbf{h}}$ and the experimental data $\mathbf{h}$, and $\mathbf{h}_{R}$ is the vector of the absolute values of the real part $\mathbf{h}$. It is used to change the weight given to the frequency $\omega_{i}$ in $\mathbf{\delta h_{R}}$. $\mathbf{\delta h_{I}}$ and $\mathbf{h}_{I}$ do the same for the imaginary part:
$\delta h_{R,i}$ $=\left|\mathrm{Re}[h_{i}-\hat{h}_{i}]\right|,$ $h_{R,i}$ $=\left|\mathrm{Re}[h_{i}]\right|+.5,$ (30)
$\delta h_{I,i}$ $=\left|\mathrm{Im}[h_{i}-\hat{h}_{i}]\right|,$ $h_{I,i}$ $=\left|\mathrm{Im}[h_{i}]\right|+.5.$ (31)
@@
Figure 4: (a) 2D array of square Ag pillars of width $w=85$ nm and height $h=35$ nm, and period $p=130$ nm over a substrate of Ag. (b) The signal is the 0-th order reflection coefficient $h(\omega)=r_{00}(\omega)$ of the 2D array depicted in (a) illuminated in normal incidence reconstructed with the SEM in Equation (26) (red curve), and compared to simulated data (blue markers). The absolute error $|h_{i}-\hat{h}_{i}|$ (black curve) is less than $10^{-3}$ at all frequencies, and a relative $L_{2}$ error of $4.32\times 10^{-3}$ % is obtained. (c) Distribution of poles in the complex $\omega$ plane.
Using the example of the 0-th order reflection coefficient of a 2D grating of silver described in Figure 4 (a), we show the result of the retrieval of the singularity expansion of Equation (26) in Figure 4 (b,c), with $M_{I}=0$ imaginary poles, $M_{C}=7$ complex pairs of poles, and $\mathbf{\alpha}=(1,0,0.2,0.2)$. The non-imaginary poles were initialized with their real parts evenly distributed from $2.5$ to $4.9\times 10^{15}$ rad/s, and their imaginary part set to 0.05-0.05 times their real parts. We obtain a relative $L_{2}$ error of $4.32\times 10^{-3}$%, with a Hermitian-symmetric solution.